6. Suppose we have N chips marked 1,2, ..., N, respectively. We take a random Hample of size 2n +1 without replacement. Let Y be the median of the random Hample. Show that the probability function of Y is for k-n +1, n +2, ..., N - n. 2n +1 Verify and Var(Y) 8n +12

Will give thumbs up for good answer!
Answer #1

These Nehis 1 2, To g medionk , s hould manked reatethan k ら 欠> n+1 Because te smallest n ho ka median and aqest median , N,Page: k (k- N-K γ) 2n N-k N-K วา างSummathon io tan teiat k+2 A-k to be kil Th nem e) E(r)= (m+ 1. 1 2n1 9 忖 nt (N- 2 t2nt2) 2Noui use shall 2n1 K = n+1 2nt 2n+ ) N-k Kent, (m+2. าว 01 mum bunsDate: Page: 2 ESY 2242 2n+2 2 2_2n2 2 202nt 3) 2